One of the common aircrafts used by air carriers from JAX to CLT is Airbus A319,...

  1. One of the common aircrafts used by air carriers from JAX to CLT is Airbus A319, which has a capacity of 128 seats. Suppose the approximate total cost (e.g., labor, fuel, landing fee, aircraft maintenance and usage, etc.) of flying JAX to CLT is $20,000.
  1. What will be unit cost (cost per passenger) when the air carrier has a load factor of 60%?

  1. What will be unit cost (cost per passenger) when the air carrier has a load factor of 85%?
  1. Compare the answers from a and b, how the change of load factor can affect the unit cost? And what strategy air carriers usually implement to maximize the load factor?
Answer #1

In order to cover the cost of flying the Airbus A319, the air carrier must have the revenue equal (at breakeven) or more than the operational cost in order to survive in the aviation business.

The given operational cost of flying Airbus A319 from JAX to CLT is = $20,000.

Given capacity of the Airbus A319 carrier = 128 passenger seats

a) Now, calculating the unit seat cost at load factor = 60%

The load factor means out of 128 seats only 60% of seats were occupied during operation and the rest of the seats went empty.

Therefore, the air carrier will have to cover the operational cost of flying Airbus A319 from JAX to CLT with this 60% of occupied seats out of 128 seats.

60% load factor = 60% x total seats = (60/100)*128 = 76.8 seats or 77 seats (As seats cannot be partially occupied)

So, in order to cover the operational cost of $20,000 with 77 occupied seats, the price of each unit of seat will be = 20,000/77 = $259.74 = $260 (approx.)

b) Now, calculating the unit seat cost at load factor = 85%

Similarly, the load factor means out of 128 seats only 85% of seats were occupied during operation and rest of the seats went empty. i.e.,

85% load factor = 85% x total seats = (85/100)*128 = 108.8 seats or 109 seats (As seats cannot be partially occupied)

Therefore, the air carrier will have to cover the operational cost of flying Airbus A319 from JAX to CLT with these 109 occupied seats out of 128 seats.

So, in order to cover the operational cost of $20,000 with 109 occupied seats, the price of each unit of seat will be = 20,000/109 = $183.48 = $184 (approx.)

c) With respect to the answers a) and b) the operational costs decrease with increase in load factor. In other words, the more people will travel the Airbus A319 from JAX to CLT, the cheaper will be the rate cost per seat for the airlines.

However, this is not the benefit for the customers, as airlines charges dynamic pricing on per seat basis (the charges of each seat will increase with the no. of seats booked). This is done to avoid selling out in advance or saving the seats for business and urgent travelers. As urgent travelers and business travelers are willing to pay more for the frequent or urgent travels, hence it is comparatively more beneficial for the airlines to collect greater share of revenues from those people.

The airlines, in order to maximize their load factors, they offer very competitive low price for the early buyers, more space and amenities to the business class travelers and sometimes they do overbook the flights in order to avoid any vacant seats. The goal is to maximize the resources and get the most profit out of each flight operations.